OpenStudy (sh3lsh):

Solve DE with undetermined coefficients y''+9y=t^2e^(3t)+6

OpenStudy (sh3lsh):

Characteristic equation solution = c1cos(3t) + c2sin(3t) Assume 2 solutions, right? What are their forms?

OpenStudy (sh3lsh):

Y1 = At^2e^(3t) Y2 = B correct? no?

OpenStudy (sh3lsh):

Not correct Y1 may equal to At^2e^(3t) + Bte^(3t) + Cte^(3t) + D

OpenStudy (sh3lsh):

Oh gosh this is too much.

OpenStudy (loser66):

oh no \(Y_1= (At^2 +Bt + C)e^{3t}\)

OpenStudy (loser66):

\(Y_1'= 3e^{3t}(At^2+Bt+C) +e^{3t}(2At+B)\) not bad :)

OpenStudy (loser66):

Now, you do hehehe. what is \(Y_1"\)?

hartnn (hartnn):

better take the constant too, otherwise we will have constant 6 on right and no constant on left :P \(Y_1= (At^2 +Bt + C)e^{3t}+ D\)

OpenStudy (sh3lsh):

Haha. @Loser66 Out of Undermined Coefficients, Laplace Transforms or Variations of Parameters, what would be the quickest method to solve this equation?

hartnn (hartnn):

Though you didn't ask me, but I am very much tempted to answer that! :P Laplace all the way!

OpenStudy (sh3lsh):

Haha, I'd love your input as well!!

OpenStudy (sh3lsh):

@hartnn :)

OpenStudy (sh3lsh):

What does one do when initial conditions aren't given?

hartnn (hartnn):

Assume them to be = 0

OpenStudy (loser66):

Whatever, we have to know all. If the question asks you to use Undetermined coefficient, do it. period

OpenStudy (sh3lsh):

Oh?

OpenStudy (sh3lsh):

(the question doesn't do Undetermined Coefficients). We have 10 minutes to solve questions like this in class. I thought Undermined would be fastest, but I'm severely regretting it!

hartnn (hartnn):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Loser66 Whatever, we have to know all. If the question asks you to use Undetermined coefficient, do it. period \(\color{#0cbb34}{\text{End of Quote}}\) 

OpenStudy (sh3lsh):

Hahaha, you all are the best.

OpenStudy (loser66):

hahahah..... I am not a boss I am a LOSER

hartnn (hartnn):

I'd say, use the method you have practiced the most with! I always preferred Laplace over other methods, and can do any ODE within few minutes using laplace, just because I have enough practice of it.

OpenStudy (loser66):

10mins? OK!! Enough

OpenStudy (sh3lsh):

Eh, I'm prone to make algebraic errors. However, partial fractions can result in errors as well.

OpenStudy (loser66):

How about Variation parameter?

hartnn (hartnn):

in undetermined co-efficient method, if your initial guess is wrong, you've wasted the time :P so you may avoid it.

OpenStudy (sh3lsh):

Is it easier? I just imagined taking that my particular solution would have the integral of trigonometric functions multiplied by that nonhomogenous term would be terrible!

OpenStudy (sh3lsh):

Ahhh so I have \[Y(s) = \frac{ \frac{ 6 }{ s } +\frac{ 4 }{ (s-3)^3 }}{ s^2 -9}\] On the right track?

OpenStudy (sh3lsh):

Err rather, s^2+9 on the denominator

hartnn (hartnn):

1. s^2 + 9 2. 2/(s-3)^3

OpenStudy (sh3lsh):

Are you sure?

hartnn (hartnn):

2! = 2

OpenStudy (sh3lsh):

Oh! Duh! Sorry!

hartnn (hartnn):

this partial fraction is terrible :O

OpenStudy (sh3lsh):

Yeah :(

hartnn (hartnn):

*Abort mission* Go back to Undetermined co-eff :P

OpenStudy (sh3lsh):

I am very confused at how anyone sane can solve this under 10 minutes.

OpenStudy (sh3lsh):

Hahaha

hartnn (hartnn):

\(Y_1= (At^2 +Bt + C)e^{3t}+ D\) Y', Y'' then compare the co-efficients on both sides

OpenStudy (sh3lsh):

How can one quickly differentiate y1?

hartnn (hartnn):

1. Practice 2. Copy from someone nearby :P

OpenStudy (sh3lsh):

Hahahahhha

OpenStudy (sh3lsh):

Will do manually

hartnn (hartnn):

wise choice :)

OpenStudy (loser66):

You can do it in 10 minutes for sure.

hartnn (hartnn):

especially since direct comparison of co-efficient gives, B= 0, C= 0, 9D = 6, 18A =1 not too bad!

hartnn (hartnn):

This makes me think, \(\color{#0cbb34}{\text{Originally Posted by}}\) @sh3lsh Y1 = At^2e^(3t) Y2 = B correct? no? \(\color{#0cbb34}{\text{End of Quote}}\) ^^ should have worked

OpenStudy (sh3lsh):

Hmmm

OpenStudy (sh3lsh):

Perhaps we're doing undermined Coefficients incorrectly (or rather, doing extraneous steps?)

hartnn (hartnn):

I thought Y = At^2 e^(3t) +B was already tried...it would have taken so much less time

OpenStudy (sh3lsh):

Is this correct: If the non homogeneous term is simply a power of t, we must try itself and the lower powers say, = t^3 Must try At^3 +Bt^2 + Ct + D If it's a power of t multiplied by exp, just try itself =t^3e^(3t) Try At^3e^(3t) (of course, multiply guess by t if it's in the characteristic solution)

OpenStudy (sh3lsh):

Or must we try the lower powers of t^3e^(3t)

hartnn (hartnn):

Yes, though i would not use the word "must" we can try whatever the hell we want! and here comes the ODE champion! :P

myininaya (myininaya):

hey how did you get B=0; I have gotten B=-1/27 I know your A is right And I haven't checked the other constants yet

hartnn (hartnn):

oh? i relied on wolfram alpha...

OpenStudy (sh3lsh):

I have to head out! Thanks all!

myininaya (myininaya):

\[y=e^{3t}(At^2+Bt+C)+D \text{ should work for the particualr solution }\]